# An Investigation of Robustness of LLMs in Mathematical Reasoning: Benchmarking with Mathematically-Equivalent Transformation of Advanced Mathematical Problems

### v1

@@ Abstract
In this paper, we introduce a systematic framework beyond conventional method to assess LLMs’ mathematical‑reasoning robustness by stress‑testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 49 % on the originals but drops by 4 percentage points on surface variants, and by 10.5 percentage points on core-step-based variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
Introduction
@@
Motivation
Modern AI systems are increasingly entrusted with tasks that hinge on reasoning rather than pattern matching. Reliable progress therefore depends on precisely measuring an LLM’s reasoning capacity and its ability to generalize beyond memorized textual surface forms. Existing math-reasoning benchmarks, however, exhibit two critical weaknesses: (i) leakage-induced score inflation, since benchmark items rapidly seep into pre-training corpora, and (ii) limited robustness coverage, because today’s datasets are too small or lack controlled transformations that probe true generalization. Addressing these weaknesses is urgent if we aim to benchmark reasoning with the same rigor demanded in safety-critical domains such as healthcare or cybersecurity.
Benchmark inflation through training leakage.
Recent studies show that public datasets, including GSM8K(Cobbe et al. 2021) and MATH (Hendrycks et al. 2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
Competition mathematics reveals the next robustness bottleneck.
@@
Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (Huang et al. 2025), and PutnamBench (640 formalized theorems) (Tsoukalas et al. 2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
Generalization–and–Perturbation (GAP).
@@
We address both leakage and robustness with a simple idea: stress‑test the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence‑preserving operators (§ 2 gives the formal definition). We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms.
PutnamGAP: instantiating GAP on 85 years of problems.
@@
We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals) and expand each item into five variants—four surface renames and one kernel rewrite—obtaining 6 306 stress‑test questions.A two‑stage QA pass—15 rounds of o3 self‑review plus a 10% spot‑check found no substantive errors.
Headline results.
@@
Across ⟨17⟩ models, as shown in 8, all of them suffer from both simple renaming and step-based rewrites. OpenAI’s o3 scores ⟨49⟩% on original statements but loses ⟨4⟩ pp under surface renames and ⟨10.5⟩ pp under kernel rewrites. These drops confirm that high leaderboard scores can collapse when cosmetic or structural perturbations are applied—precisely the effect that data leakage masks.
Contributions.
@@
(1) We introduce GAP, a general framework for measuring robustness via mathematically equivalent transformations. (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across seventeen LLMs, plus an open‑source evaluation stack.
Method
We start from a curated set of $N$ canonical items
$\mathcal{P}=\bigl{\{}(x_{i},y_{i},\pi_{i})\bigr{\}}_{i=1}^{N},$
@@
where $x_{i}$ is a problem statement, $y_{i}$ is its reference answer(s), and $\pi_{i}$ an unreleased expert solution path used internally for safe variant generation.
Model interface.
A language model $f_{\theta}$ receives a prompt $x$ and returns $\hat{y}=f_{\theta}(x)$, which an automatic checker maps to a binary label $z=\operatorname{grade}(\hat{y},y)\in\{0,1\}$.
Variant families.
For every $x_{i}$ we later apply two disjoint transformation super‑families (defined in the next section but left unchanged here):
$\mathcal{T}_{i}^{\textsf{surf}}$ $\;(\text{$K_{surf}$ surface variants}),$ (1)
$\mathcal{T}_{i}^{\textsf{para}}$ $\;(\text{$K_{core}$ parametric variant}).$ (2)
@@
Each surface transformation $\tau$ returns a new statement $x_{i}^{(\tau)}=\tau(x_{i})$ that preserves semantic correctness of $y_{i}$. For parametric variations, $y_{i}$ is transformed as well to match $\tau(x_{i})$
Evaluation matrix.
@@
The Cartesian product $\mathcal{D}=\{(i,\tau)\mid i\leq N,\;\tau\in\mathcal{T}_{i}^{\textsf{surf}}\cup\mathcal{T}_{i}^{\textsf{para}}\cup\{\mathrm{id}\}\}$ contains $N\times(K+1)$ aligned items (original + $K$ variants per source, $K=K_{surf}+K_{para}$). Running $f_{\theta}$ on every pair populates a binary matrix $\mathbf{Z}\in\{0,1\}^{N\times(K+1)}.$ From the first column we extract the easy vector $\mathbf{e}(\theta)\in\{0,1\}^{N},$ while the remaining columns feed family‑specific aggregates:
$\mathbf{h}^{\textsf{surf}}(\theta)=\text{maj}\bigl{(}\mathbf{Z}_{[:,\,\textsf{surf}]}\bigr{)},\quad\mathbf{h}^{\textsf{para}}(\theta)=\mathbf{Z}_{[:,\,\textsf{para}]}\!.$
The set of surface variants can be changed based on specific tasks.
Robustness Metric
Let $\mathbf{e},\mathbf{h}\!\in\!\{0,1\}^{N}$ be per-item correctness on the easy and hard sets. In this work, easy sets refer to the original and hard sets refer to the variants. Let $p_{e},p_{h}$ be their Jeffreys-smoothed accuracies, and define the pooled standard deviation $\sigma=\sqrt{\tfrac{1}{2}\bigl{[}p_{e}(1-p_{e})+p_{h}(1-p_{h})\bigr{]}}$. We measure robustness with
$R(\mathbf{e},\mathbf{h})=\frac{1}{N}\sum_{j=1}^{N}\exp\!\Bigl{(}-\tfrac{e_{j}-h_{j}}{\sigma}\Bigr{)}.$
The exponent converts accuracy drops (in $\sigma$ units) into a multiplicative penalty in $(0,1]$: $R=1$ means perfect invariance, while smaller values indicate greater brittleness. Full derivation, statistical justification, and design discussion are in Appendix B.
Robustness scores.
Using the robustness metric, we report
$R_{\textsf{surf}}$ $=R\bigl{(}\mathbf{e},\mathbf{h}^{\textsf{surf}}\bigr{)},\quad R_{\textsf{core}}=R\bigl{(}\mathbf{e},\mathbf{h}^{\textsf{para}}\bigr{)},$
$R_{\textsf{global}}$ $=\sqrt{R_{\textsf{surf}}\,R_{\textsf{para}}}.$ (3)
Only these vectors are needed—the content of the Transformation Families subsection itself is left intact.
Reproducibility.
An open‑sourced script will be released after review.
Transformation Families
@@
The proposed general robustness measures can work for any variations. As a first step in exploring this new evaluation methodology, we propose and study five aligned variants— four surface renamings that perturb only symbol names, and one core‑step instance that perturbs numeric slots while preserving the reasoning chain. This section details the synthesis pipelines;
Surface renaming family
@@
    Collision check. A deterministic post‑validator rejects names colliding with any pre‑existing identifier in the problem.
@@
    Family tagging. The accepted string is labelled as belonging to one of four families described below.
The Five families
DL
Single descriptive phrase (populationDensity).
DLC
2–5 unrelated nouns (walnutVioletTerrace).
DLM
Misleading mathematical term (primeOrder).
GS
4–16‑char hash (xcQ7h2ZfRw9v).
@@
Parametric variant
Symbol renaming probes only the lexical axis. To probe structural transfer, we resample numerical constants yet force the solution to reuse the original high‑level moves. In this work, we call it Kernel_Variant (KV).
Four‑stage pipeline
We convert each item into semantically-equivalent variants through a four-stage pipeline (pseudocode (Alg. 1)):
  1. 1.
    slot discovery,
  2. 2.
    template back-synthesis,
  3. 3.
    semantic-preserving renaming, and
  4. 4.
    dual-verifier screening (two-in-a-row rule).
The pipeline generates a bounded number of validated variants for each problem within a few hours on commodity hardware using the OpenAI o3 API. See Appendix C for empirical bounds and implementation details.
Algorithm 1 Repair‑and‑verify loop (excerpt)
1:input: draft variant $v_{0}$
2:for $t=1$ to $T$ do
3:  Run $J$ o3 judges $\to$ verdict vector $\mathbf{z}_{t}$

### v2

@@ Abstract
In this paper, we introduce a systematic framework beyond conventional method to assess LLMs’ mathematical-reasoning robustness by stress-testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 51.5 % on the originals but drops by 4.7 percentage points on surface-renaming variants, and by 12.9 percentage points on parametric variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
1  Introduction
Motivation. Modern AI systems are increasingly entrusted with tasks that hinge on robust reasoning rather than pattern matching. It is thus important to precisely measure an LLM’s reasoning capacity and its ability to generalize beyond memorized textual surface forms. Existing math-reasoning benchmarks, however, exhibit two critical weaknesses: (i) leakage-induced score inflation, since benchmark items rapidly seep into pre-training corpora, and (ii) limited robustness coverage, because today’s datasets are too small or lack controlled transformations that probe true generalization. Addressing these weaknesses is urgent if we aim to benchmark reasoning with the same rigor demanded in safety-critical domains such as healthcare or cybersecurity.
Benchmark inflation through training leakage. Recent studies show that public datasets, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
@@ 1  Introduction
Competition mathematics reveals the next robustness bottleneck. Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (Huang et al., 2025), and PutnamBench (640 formalized theorems) (Tsoukalas et al., 2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
@@
Generalization–and–Perturbation (GAP), a novel evaluation strategy. We address both leakage and robustness with a new idea: stress‑test the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence‑preserving operators. We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms. It offers a new general diagnostic evaluation methodology for analyzing and quantifying the robustness of an LLM’s mathematical reasoning capacity.
@@
PutnamGAP: instantiating GAP on 85 years of problems. We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1,051 originals) and expand each item into five variants—four surface renames and one kernel rewrite—obtaining 6,306 stress‑test questions. A two‑stage QA pass—15 rounds of o3 self‑review plus a 10% spot‑check found no substantive errors.
@@
Headline results. Across 18 models, as shown figure 4, all of them suffer from both simple renaming and step-based rewrites. OpenAI’s o3 scores 51.5% on original statements but loses 4.7 pp (9.12%) under surface renames and 12.9 pp (25.22%) under parametric rewrites. These drops confirm that high leaderboard scores can collapse when cosmetic or structural perturbations are applied—precisely the effect that data leakage masks.
@@
Contributions. (1) We propose GAP, a novel general framework for measuring robustness via mathematically equivalent transformations that overcomes two common deficiencies of the current evaluation methods (i.e., data leakage and lack of robustness measures). (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across seventeen LLMs, plus an open‑source evaluation stack.
2  The Generalization–and–Perturbation (GAP) Framework
2.1 Evaluation Model
@@
We start from a curated set of $N$ canonical items $\mathcal{P}=\{(x_{i},y_{i},\pi_{i})\}_{i=1}^{N}$, where $x_{i}$ is a problem statement, $y_{i}$ is its reference answer(s), and $\pi_{i}$ an unreleased expert solution path used internally for safe variant generation. Model interface. A language model $f_{\theta}$ receives a prompt $x$ and returns $\hat{y}=f_{\theta}(x)$, which an automatic checker maps to a binary label $z=\operatorname{grade}(\hat{y},y)\in\{0,1\}$.
@@
Variant families. For every $x_{i}$ we later apply two disjoint transformation super-families (defined in the next section but left unchanged here): $\mathcal{T}_{i}^{\textsf{surf}}$ ($K_{surf}$ surface variants), $\mathcal{T}_{i}^{\textsf{para}}$ ($K_{para}$ parametric variants). Each surface transformation $\tau$ returns a new statement $x_{i}^{(\tau)}=\tau(x_{i})$ that preserves semantic correctness of $y_{i}$. For parametric variations, $y_{i}$ is transformed as well to match $\tau(x_{i})$.
@@
Evaluation matrix. The Cartesian product $\mathcal{D}=\{(i,\tau)\mid i\leq N,\;\tau\in\mathcal{T}_{i}^{\textsf{surf}}\cup\mathcal{T}_{i}^{\textsf{para}}\cup\{\mathrm{id}\}\}$ contains $N\times(K+1)$ aligned items (original + $K$ variants per source, $K=K_{surf}+K_{para}$). Running $f_{\theta}$ on every pair populates a binary matrix $\mathbf{Z}\in\{0,1\}^{N\times(K+1)}$. From the first column we extract the easy vector $\mathbf{e}(\theta)\in\{0,1\}^{N}$, while the remaining columns feed family-specific aggregates: $\mathbf{h}^{\textsf{surf}}(\theta)=\text{maj}(\mathbf{Z}_{[:,\,\textsf{surf}]})$, $\mathbf{h}^{\textsf{para}}(\theta)=\mathbf{Z}_{[:,\,\textsf{para}]}\!$. The set of surface variants can be changed based on specific tasks.
Robustness Metric. Let $e,h\in\{0,1\}^{N}$ denote per-item correctness on the easy (original) and hard (variant) sets. With Jeffreys smoothing
$p_{e}=\frac{\sum_{j}e_{j}+\tfrac{1}{2}}{N+1},\quad p_{h}=\frac{\sum_{j}h_{j}+\tfrac{1}{2}}{N+1},\quad\sigma=\sqrt{\tfrac{1}{2}\big(p_{e}(1-p_{e})+p_{h}(1-p_{h})\big)}.$
Define the SD-normalized drop $d_{j}=(e_{j}-h_{j})/\sigma$ and its soft-saturated version $\widehat{d}_{j}=\tfrac{1}{k}\log\!\big(1+e^{kd_{j}}\big)$ with $k\approx 0.5$. Let $\tilde{d}=\operatorname{median}\{d_{j}\,|\,d_{j}>0\}$ (with fallback $\tilde{d}:=\max(\varepsilon,\operatorname{median}|d_{j}|)$, $\varepsilon=0.1$ when no positive drop exists) and set $\beta=\ln 2/\tilde{d}$. Our penalty robustness is
$\widehat{R}(e,h)=\frac{1}{N}\sum_{j=1}^{N}\exp\!\big(-\beta\,\widehat{d}_{j}\big)\in(0,1].$
Thus $\widehat{R}=1$ indicates invariance; a “typical” loss ($\widehat{d}_{j}\approx\tilde{d}$) halves the per-item factor, while improvements ($d_{j}<0$) are clamped to zero penalty (no reward). We report $R_{\text{surf}}=\widehat{R}(e,h_{\text{surf}})$, $R_{\text{para}}=\widehat{R}(e,h_{\text{para}})$, and $R_{\text{global}}=\sqrt{R_{\text{surf}}R_{\text{para}}}$. Full derivation, statistical justification, and design discussion are in Appendix B.
2.2 Transformation Families
@@
The proposed general robustness measures can work for any variations. As a first step in exploring this new evaluation methodology, we propose and study five aligned variants— four surface renamings that perturb only symbol names, and one core‑step instance that perturbs numeric slots while preserving the reasoning chain. This section details the synthesis pipelines. Detailed descriptions can also be found in Appendix A.
2.2.1 Surface renaming variant family
Figure 2: Surface renaming variant family pipeline
@@ 2.2.1 Surface renaming variant family
    Collision check. A deterministic post-validator rejects names colliding with any pre-existing identifier in the problem.
@@
    Family tagging. The string is labelled as belonging to one of four families described below.
We use four types of surface variants: Descriptive_Long (DL), with a single descriptive phrase; Descriptive_Long_Confusing (DLC), with 2–5 random unrelated nouns; Descriptive_Long_Misleading (DLM), with a mathematically suggestive but misleading term ; Garbled_String (GS), with a 4–16-character hash, as shown in figure 2 where ‘Q’ stands for the problem question and ‘A’ stands for the official solution.
@@
2.2.2 Parametric variant family
Figure 3: Parametric variant family pipeline
Symbol renaming probes only the lexical axis. To probe structural transfer, we resample numerical constants yet force the solution to reuse the original high-level moves. In this work, we call it Kernel_Variant (KV). We convert each item into semantically-equivalent variants through a four-stage pipeline: (1) slot discovery; (2) template back-synthesis; (3) question reverse-engineering; and (4) dual-verifier screening (two-in-a-row rule). The pipeline generates a bounded number of validated variants for each problem within a few hours on commodity hardware using the OpenAI o3 API. See Appendix A for empirical bounds and details of our implementation.
2.3 Implementation Overview
Code release. To facilitate double‑blind reviewing we publish only the subset of data (100 randomly chosen examples). An automated evaluator, putnam-cli.py, receives the names of target solver model and grader model and variant type to test. Supported back‑ends are (i) any HuggingFace‐compatible checkpoint via transformers, (ii) a local vllm server, or (iii) API clients including OpenAI, Gemini, Anthropic and OpenRouter. Full data and generation scripts will be released post‑decision.
Surface generation. Renaming variants are produced on a CPU‑only node by streaming o3 API calls. A five‑stage exponential‑back‑off retry (max 5 attempts, doubling timeout each time) masks transient API latency. Processing all 1 0511\,051 items in parallel takes $\sim$15 min wall‑clock.
Core‑step generation. Kernel variant synthesis is more expensive because of multi‑turn chain‑of‑thought reasoning: end‑to‑end runtime is $\leq\!3\,\text{h}$ for the full corpus on a single 8‑core CPU, dominated by the 15‑iteration repair‑and‑verify loop.
3  PutnamGAP Dataset
3.1 Data Sources, Extraction & Annotation
Our benchmark comprises all Putnam Problems 1938–2024 (N=1 051N=1\,051 items after deduplication). See Appendix E for source details.
Original scans are processed via a 3-stage OCR routine: (i) Manual segmentation for every question-answer pair. (ii) MathPix for formula‐aware PDF-to-LaTeX conversion followed by (iii) custom post-filters that merge multi-line expressions and fix 4̃.2 % residual symbol errors. Each item is manually spot-checked ($\leq$2 min per problem) to ensure semantic fidelity before variant generation. Complete corpus list, OCR accuracy study, and cleaning scripts appear in Appendix E.
3.2 Dataset Statistics
Overall scale and balance. The benchmark comprises 1,051 original Putnam problems from 1938–2024 and five mathematically equivalent transformations, yielding 6,306 items. Part distribution is balanced (527 A vs. 524 B), and the canonical identifier $\langle\textit{year},\textit{part}\{A,B\},\textit{index}\rangle$ provides a difficulty proxy. Using indices 121\!\!-\!2 as Easy, 343\!\!-\!4 as Medium, and 565\!\!-\!6 as Hard, the corpus contains 32.3 % Easy, 32.3 % Medium, 32.2 % Hard, plus a 3.0 % extra–hard tail (indices 7–8).
Topic coverage and Quality Control Automatic tags in _meta.tag indicate broad mathematical coverage—Algebra (641), Analysis (521), Number Theory (392), Combinatorics (286), and Geometry (239). 803 of the questions are proofs, and 248 of them are calculations. At the same time, every item has undergone single‐pass manual validation.
4  Experimental Setup
The constructed PutnamGAP dataset enables, for the first time, a robust analysis of an LLM’s reasoning capacity. In this section, we describe how we set up the experiments to evaluate the robustness of 18 representative models.
4.1 Model Pool & Prompting
We evaluated 18 models (see 1 or Appendix A for a complete list).All models are queried under a unified zero‑shot template. A system instruction designates the model as “an expert mathematician” and asks it to show all work, while the user message embeds the problem. See Appendix G for our full prompt. We fix temperature=0, top_p=1, and max_tokens=32000 or maximum token amount available in case some models have max_tokens maximum smaller than 32000. for every run except OpenAI O-series which require temperature=1. Solutions are then re‑submitted to a second template that grades the answer: a strict proof rubric for proof items and a lenient numeric rubric for calculation items. Both grader prompts require structured JSON output containing a binary grade field plus detailed feedback. Complete prompt code is available in Appendix G
4.2 Scoring & Auto-Grader
We partition tasks into computation and proof categories and evaluate them with distinct graders.
Computation Each candidate answer is normalized (whitespace, units, LaTeX macros) and passed to two scoring paths: (i) a strict string match against the reference solution; (ii) a latent grader—an LLM prompted to return ‘‘CORRECT’’ or ‘‘INCORRECT’’ given the reference answer and a rubric that disallows partial credit. We adopt path (ii) to mitigate formatting artifacts; if the two paths disagree we mark the item for manual audit (¡1% of cases).
Proof We provide the grader with an aligned, step-by-step reference proof and ask it to assign a binary grade plus a natural-language justification. Any skipped logical step or missing citation triggers a fail. A random 10 % sample is double-checked by independent volunteers; grader precision/recall is >>97 %.
5  Results & Analysis
5.1 Robustness
Table 1: Model Accuracy Rates across Categories (Percent Scale)
Model DL ($\Delta$) DLC($\Delta$) DLM ($\Delta$) GS ($\Delta$) Kernel Variant ($\Delta$)
claude-opus-4 23.0∗∗ (–3.5) 22.2∗∗∗ (–4.3) 21.7∗∗∗ (–4.8) 21.4∗∗∗ (–5.1) 13.8∗∗∗ (–12.7)
claude-sonnet-4 20.6∗∗ (–2.5) 19.8∗∗∗ (–3.2) 18.6∗∗∗ (–4.4) 18.1∗∗∗ (–4.9) 11.1∗∗∗ (–11.9)
deepseek-prover 15.2∗∗∗ (–0.3) 14.0∗∗∗ (–1.5) 12.8∗∗ (–2.7) 13.7∗∗ (–1.8) 9.2 ∗∗∗ (–6.3)
gemini-2.5-flash-lite 18.8∗∗∗ (–0.9) 16.1∗∗∗ (–3.7) 15.8∗∗∗ (–4.0) 15.1∗∗∗ (–4.7) 6.6 ∗∗∗ (–13.2)
gemini-2.5-pro 75.2∗∗ (–3.1) 74.3∗∗∗ (–4.1) 72.8∗∗∗ (–5.6) 72.9∗∗∗ (–5.4) 63.5∗∗∗ (–14.9)
gemini-2.5-flash 42.6∗∗∗ (–0.2) 39.0∗∗∗ (–3.8) 40.9∗∗∗ (–1.9) 37.6∗∗∗ (–5.2) 27.6∗∗∗ (–15.2)
gpt-4.1 23.4∗∗∗ (–1.5) 21.2∗∗ (–3.7) 22.0∗∗ (–2.9) 22.6∗∗∗ (–2.3) 14.8∗∗∗ (–10.1)
gpt-4.1-mini 26.9∗∗ (–1.7) 27.1∗∗∗ (–1.5) 24.0∗∗∗ (–4.6) 25.1∗∗ (–3.4) 19.2∗∗∗ (–9.4)
gpt-4.1-nano 8.9 ∗∗∗ (+0.1) 6.4 ∗∗ (–2.4) 7.3 ∗∗ (–1.5) 6.6 ∗∗ (–2.2) 6.8 ∗∗∗ (–2.0)
gpt-4o 6.3 ∗∗∗ (–0.1) 5.3 ∗∗ (–1.1) 4.7 ∗∗ (–1.8) 4.7 ∗∗∗ (–1.8) 3.0 ∗∗∗ (–3.4)
gpt-4o-mini 3.5 ∗∗∗ (–0.8) 2.5 ∗∗∗ (–1.8) 3.3 ∗∗∗ (–1.0) 3.2 ∗∗∗ (–1.1) 1.7 ∗∗∗ (–2.6)
grok4 59.0∗∗∗ (–1.1) 56.6∗∗∗ (–3.4) 55.9∗∗∗ (–4.2) 55.2∗∗∗ (–4.8) 45.5∗∗∗ (–14.6)
kimi-k2 25.8∗∗∗ (–1.4) 23.7∗∗ (–3.5) 23.8∗∗ (–3.4) 23.4∗∗∗ (–3.8) 12.8∗∗∗ (–14.4)
llama-4 14.5∗∗∗ (–1.2) 13.0∗∗ (–2.7) 13.1∗∗ (–2.6) 13.8∗∗ (–1.9) 7.3 ∗∗∗ (–8.4)
mistral 5.5 ∗∗∗ (–0.1) 5.7 ∗∗∗ (+0.1) 4.9 ∗∗∗ (–0.7) 4.2 ∗∗ (–1.3) 3.9 ∗∗ (–1.6)
o3 52.8∗∗∗ (+1.3) 47.6∗∗ (–3.8) 49.5∗∗∗ (–2.0) 46.8∗∗∗ (–4.7) 38.6∗∗∗ (–12.9)
o4-mini 43.0∗∗∗ (+1.5) 38.3∗∗ (–3.2) 38.5∗∗∗ (–3.0) 40.4∗∗∗ (–1.1) 29.1∗∗∗ (–12.4)
qwen3 29.3∗∗∗ (+1.1) 27.3∗∗∗ (–0.9) 26.9∗∗∗ (–1.4) 26.5∗∗∗ (–1.7) 14.9∗∗∗ (–13.4)
Note: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we used a paired design (McNemar’s exact test) on matched problem pairs to test whether the accuracy rate decreases significantly compared to the original. Statistically significant differences are indicated using standard notation ($p<0.1$, $p<0.05$, $p<0.01$). We also computed 95 % CI (See Appendix D Figure 4 ) and robustness metrics $R$ (See Appendix D Figure 9 ), and all models, especially those performed well on the original set, got a low robustness score.
Table 2: Robustness metrics $R_{\text{surf}},R_{\text{para}},R_{\text{global}}$ (rounded to three decimals).
Model $R_{\text{surf}}$ $R_{\text{para}}$ $R_{\text{global}}$ Model $R_{\text{surf}}$ $R_{\text{para}}$ $R_{\text{global}}$
claude-opus-4 0.958 2360.958\,236 0.948 9230.948\,923 0.953 5690.953\,569 gpt-4o 0.985 8360.985\,836 0.980 0260.980\,026 0.982 9270.982\,927
claude-sonnet-4 0.961 3280.961\,328 0.941 5450.941\,545 0.951 3850.951\,385 gpt-4o-mini 0.989 6640.989\,664 0.985 6520.985\,652 0.987 6560.987\,656
deepseek-prover 0.972 3610.972\,361 0.959 8140.959\,814 0.966 0670.966\,067 grok4 0.937 4150.937\,415 0.916 0480.916\,048 0.926 6700.926\,670
gemini-2.5-flash-lite 0.961 2020.961\,202 0.942 2020.942\,202 0.951 6550.951\,655 kimi-k2 0.955 1850.955\,185 0.929 6170.929\,617 0.942 3140.942\,314
gemini-2.5-pro 0.949 1320.949\,132 0.914 8770.914\,877 0.931 8480.931\,848 llama-4 0.972 0190.972\,019 0.954 6040.954\,604 0.963 2720.963\,272
gemini_2.5_flash 0.951 5780.951\,578 0.917 5230.917\,523 0.934 3960.934\,396 mistral 0.984 0810.984\,081 0.981 9830.981\,983 0.983 0320.983\,032
gpt-4.1 0.963 0350.963\,035 0.944 3500.944\,350 0.953 6460.953\,646 o3 0.939 7310.939\,731 0.921 0620.921\,062 0.930 3500.930\,350
gpt-4.1-mini 0.953 1850.953\,185 0.938 5900.938\,590 0.945 8600.945\,860 o4-mini 0.946 0860.946\,086 0.928 9190.928\,919 0.937 4630.937\,463
gpt-4.1-nano 0.980 2450.980\,245 0.982 3470.982\,347 0.981 2950.981\,295 qwen3 0.941 2080.941\,208 0.927 5470.927\,547 0.934 3530.934\,353